# RNA-FrameFlow: Flow Matching for de novo 3D RNA Backbone Design

### v3

@@ A.3 RhoFold Length Bias
Figure 12: RhoFold length bias. The blue dotted line represents the median RMSD of RhoFold predictions to the RNAsolo samples.
@@ A.4 Imputing Non-frame Atoms from Torsion Angles
    Compute a vector perpendicular to $\vec{u}_{AB}$ by choosing a random normal vector $\vec{n}$ (like $[1,0,0]^{T}$ or $[0,1,0]^{T}$) and taking their cross product to get $\vec{u}_{1}=\vec{u}_{AB}\times\vec{n}$.
@@
    $(x_{C},y_{C},z_{C})=(x_{B},y_{B},z_{B})+r_{BC}\cdot\vec{u}_{\text{rot}}\leavevmode\nobreak\ .$
We use predetermined bond lengths between atoms in the idealized geometry of the Adenine (A) nucleotide from OpenComplex (Jingcheng et al., 2022) in the same way Yim et al. (2023b; a) use Alanine for generated protein backbones. We use the following atom triplets and predicted torsion angles to build the all-atom nucleotide, starting from the ribose sugar ring towards the $5^{\prime}$ end:
Fixed bond Non-frame atom Torsion angle
@@ B.1 Composition of Backbone Coordinate Loss
We also analyze how changing the composition of atoms in the inter-atom losses affects performance. We increase the number of atoms being supervised in the $\mathcal{L}_{\text{bb}}$ loss described above. Aside from the frame comprising $C3^{\prime}$, $C4^{\prime}$, and $O4^{\prime}$, we try two settings with 3 and 7 additional non-frame atoms included in the loss. For the 3 non-frame atoms, we additionally choose $C1^{\prime}$, $P$, and $O3^{\prime}$, and for the 7 non-frame atoms, we choose a superset $C1^{\prime}$, $P$, $O3^{\prime}$, $C5^{\prime}$, $OP1$, $OP2$, and $N1/N9$. We posit the additional supervision may increase the local structural realism, which may further improve validity, as shown in Table 5.
@@
Frame composition in $\mathcal{L}_{\text{bb}}$ Validity $\uparrow$ Diversity $\uparrow$ Novelty $\downarrow$
@@ B.3 Choice of Forward-folding Model
In our work, we rely on RhoFold (Shen et al., 2022) to forward fold the inverse-folded sequences from gRNAde. Here, we reperform our evaluation from Section 4.1 with Chai-1 (Boitreaud et al., 2024), a recent open-source structure prediction model with results similar to AlphaFold2, replacing RhoFold in the self-consistency pipeline in Figure 2. We do not use MSAs for Chai-1.
We do not observe any major differences in self-consistency between the two forward-folding models. For RNA-FrameFlow with RhoFold, we report a validity of 41.0% while RNA-FrameFlow with Chai-1 gives a validity of 39.5%. Recent benchmarks (Tarafder et al., 2024) also observe that existing RNA structure prediction tools like RhoFold, RF2NA (Baek et al., 2022b), and trRosettaRNA (Wang et al., 2023) perform similarly due to similarities in their architectures and training data. We compare the scTM distribution among valid samples between RhoFold and Chai-1 predictions in single-sequence mode in Figure 14.
Figure 14: Correlation between RhoFold and Chai-1 scTM scores. Horizontal and vertical dotted lines indicate the average scTM score from each method among valid samples.
@@ C.4 Measuring All-atom Steric Clashes
$\mathbb{I}_{ij}$ =$=\begin{cases}1&r_{ij}\leq d_{\text{steric}}\\ 0&\text{otherwise}\end{cases}$ (12)
@@
Figure 20: All-atom steric clashes by sequence length. We observe a similar number of steric clashes between training and generated backbones across sequence lengths. We include the (validity) for generated samples from each sequence length below the length labels along the horizontal axis.

### v4

@@ A.3 RhoFold Length Bias
Figure 12: RhoFold length bias. The blue dotted line represents the median RMSD of RhoFold predictions to the RNAsolo samples. RhoFold performs well for over-represented sequence lengths in the PDB, and poorly for under-represented sequence lengths.
@@ A.4 Imputing Non-frame Atoms from Torsion Angles
    Compute a vector perpendicular to $\vec{u}_{AB}$ by choosing a random normal vector $\vec{n}$ (like $1,0,0]^{T}$ or $0,1,0]^{T}$) and taking their cross product to get $\vec{u}_{1}=\vec{u}_{AB}\times\vec{n}$.
@@
    $x_{C},y_{C},z_{C}]=[x_{B},y_{B},z_{B}]+r_{BC}\cdot\vec{u}_{\text{rot}}\leavevmode\nobreak\ .$
We use predetermined bond lengths between atoms in the idealized geometry of the Adenine (A) nucleotide from OpenComplex (Jingcheng et al., 2022) in the same way Yim et al. (2023b; a) use Alanine for generated protein backbones. We use the following atom triplets and predicted torsion angles to build the all-atom nucleotide, starting from the ribose sugar ring towards the $5^{\prime}$ end (i.e., the phosphate group):
Fixed bond
Non-frame atom
Torsion angle
@@ B.1 Composition of Backbone Coordinate Loss
We also analyze how changing the composition of atoms in the inter-atom losses affects performance. We increase the number of atoms being supervised in the $\mathcal{L}_{\text{bb}}$ loss described above. Aside from the frame comprising $\{C3^{\prime}$, $C4^{\prime}$, $O4^{\prime}\}$, we try two settings with 3 and 7 additional non-frame atoms included in the loss. For the 3 non-frame atoms, we additionally choose $\{C1^{\prime}$, $P$, $O3^{\prime}\}$, and for the 7 non-frame atoms, we choose a superset $\{C1^{\prime}$, $P$, $O3^{\prime}$, $C5^{\prime}$, $OP1$, $OP2$, N1/N9}N1/N9\}. We posit the additional supervision may increase the local structural realism, which may further improve validity, as shown in Table 5.
@@
Frame composition in $\mathcal{L}_{\text{bb}}$ % Validity $\uparrow$ Diversity $\uparrow$ Novelty $\downarrow$
@@ B.3 Choice of Forward-folding Model
Figure 14: Correlation between RhoFold and Chai-1 scTM scores. Horizontal and vertical dotted lines denote the median scTM score from each method across all samples.
[…]
Recent benchmarks (Tarafder et al., 2024) also observe that existing RNA structure prediction tools like RhoFold, RF2NA (Baek et al., 2022b), and trRosettaRNA (Wang et al., 2023) perform similarly due to similarities in their architectures and training data. For 600 generated backbones from RNA-FrameFlow, we get 8 predicted sequences from gRNAde, giving us 4800 predicted backbones from RhoFold and Chai-1. We compare this scTM distribution across predicted structures in Figure 14.
B.4 Rotational and All-atom Self-consistency
Our self-consistency pipeline currently computes TM-score and RMSD with $C4^{\prime}$ coarse graining, reflecting performance only along the translational component. However, our frames not only comprise the $C4^{\prime}$ atom but the $C3^{\prime}$ and $O4^{\prime}$ atoms as well, forming a rotational component. Factoring in this rotational component into the self-consistency metrics would offer a clearer picture of RNA-FrameFlow’s ability to precisely orient frames. In Table 7, we report scRMSD statistics over our generated backbones. We observe that the scRMSD values from RNA-FrameFlow samples correlate positively with the self-consistency scores in Figure 3 (left). For sequence lengths with high % validity, we see relatively lower scRMSD values and higher scTM scores.
Sequence length Median scRMSD $\downarrow$ Mean scRMSD $\downarrow$ Std. Dev. scRMSD $\downarrow$
40 3.81 6.01 4.15
50 3.66 9.36 10.6
60 5.12 9.36 9.98
70 3.74 6.88 8.85
80 6.06 8.43 5.95
90 10.38 11.71 6.74
100 13.27 13.54 6.65
110 14.36 13.33 6.43
120 1.89 3.034 3.89
130 17.45 16.23 5.82
140 11.17 13.06 4.90
150 20.28 20.29 4.67
Table 7: Statistics of rotational scRMSD across sequence lengths. We observe RNA-FrameFlow orients frames realistically, correlating positively with reported scTM in Section 4.1 across sequence lengths.
RNA-FrameFlow generates backbones at an all-backbone-atom granularity. We additionally compute scRMSD across all 13 nucleotide atoms in Table 8. We similarly observe RNA-FrameFlow can generate realistic fine-grained nucleotides: for sequence lengths with relatively higher % validity, we see lower all-atom scRMSD values.
Sequence length Median scRMSD $\downarrow$ Mean scRMSD $\downarrow$ Std. Dev. scRMSD $\downarrow$
40 4.22 6.36 3.88
50 4.05 9.73 10.48
60 5.41 9.65 9.85
70 4.19 7.27 8.71
80 6.20 8.70 5.78
90 10.53 11.87 6.65
100 13.26 13.65 6.57
110 14.39 13.48 6.25
120 2.82 3.85 3.67
130 17.47 16.31 5.74
140 11.29 13.19 4.85
150 20.29 20.32 4.64
Table 8: Statistics of all-backbone-atom scRMSD across sequence lengths. We observe RNA-FrameFlow generates realistic fine-grained nucleotides, correlating positively with reported EMD scores for local structural descriptors in Table 3 across sequence lengths.
@@ C.4 Measuring All-atom Steric Clashes
$\mathbb{I}_{ij}$ $=\begin{cases}1&r_{ij}\leq d_{\text{steric}}\\ 0&\text{otherwise}\end{cases}$ (12)
@@
Figure 20: All-atom steric clashes by sequence length. We observe a similar number of steric clashes between training and generated backbones across sequence lengths. We include the (% validity) for generated samples from each sequence length below the length labels along the horizontal axis.
C.5 Atomic Displacement of Frame Atoms
To further motivate our choice of frame atoms, we examine the B-factor of each atom provided in the RNAsolo PDB files. B-factor is a measure of atomic displacement, i.e., how much an atom wiggles in its place when the structure is undergoing X-ray crystallography or Cryo-EM. Empirically, we observe our frame construction $\{C4^{\prime},C3^{\prime},O4^{\prime}\}$ collectively deviates the least compared to the frame construction in RF2NA (Baek et al., 2022b), $\{P,OP1,OP2\}$. We report this in Table 9 below:
Atom Mean B-factor2) $\downarrow$ Median B-factor2) $\downarrow$
C4’ 111.44 84.84
C3’ 111.74 85.41
O4’ 111.22 84.86
P 114.03 88.69
OP1 113.46 87.56
OP2 112.35 86.11
Table 9: Statistics of B-factor values (atomic displacement) of nucleotides from 500 random samples in RNAsolo. We observe our frame construction $\{C4^{\prime},C3^{\prime},O4^{\prime}\}$ experiences lower collective spatial uncertainty compared to alternate frame constructions such as $\{P,OP1,OP2\}$.
C.6 Modeling Ring Puckering
Puckering refers to non-planar deformations or contortions of the ribose sugar ring in RNA comprising the atoms $\{C1^{\prime},C2^{\prime},C3^{\prime},C4^{\prime},O4^{\prime}\}$. Instead of forming an ideal flat plane, the ring atoms alternately move above and below the ideal plane to relieve steric strain (see Figure 21 (A)). The five atoms collectively define five out-of-plane torsion angles: $\nu_{0}$ ($C4^{\prime}-O4^{\prime}-C1^{\prime}-C2^{\prime}$), $\nu_{1}$ ($O4^{\prime}-C1^{\prime}-C2^{\prime}-C3^{\prime}$), $\nu_{2}$ ($C1^{\prime}-C2^{\prime}-C3^{\prime}-C4^{\prime}$), $\nu_{3}$ ($C2^{\prime}-C3^{\prime}-C4^{\prime}-O4^{\prime}$) and $\nu_{4}$ ($C3^{\prime}-C4^{\prime}-O4^{\prime}-C1^{\prime}$). For a nucleotide along the backbone, these angles define Cremer–Pople pseudo-rotations parameterized by a phase angle $P$, that locates which atom is displaced towards the endo position (i.e., above the flat plane), and an amplitude $\tau_{m}$, that quantifies the maximum magnitude of displacement. Tracking these quantities provides insights into the conformational diversity and interactions of RNA. Specifically, $C2^{\prime}$-endo and $C3^{\prime}$-endo puckering are known to mediate these behaviours. Several riboswitches, ribozymes, and RNA-protein interfaces exploit such endo orientations of ribose atoms as a trigger mechanism during molecular binding (Setlik et al., 1995; Salter et al., 2006).
In naturally-ocurring RNA (e.g., RNAsolo), the phase angle $P$ distribution is sharply bimodal with a dominant $C3^{\prime}$-endo near $0^{\circ}$ and a smaller $C2^{\prime}$-endo near $180^{\circ}$. The puckering amplitude $\tau_{m}$ forms a narrow band near $32^{\circ}-38^{\circ}$ (0.55-0.66 rad) (Shi et al., 2020; Harp et al., 2022). In Figure 21 (B), we demonstrate that RNA-FrameFlow captures key ring puckering motions, with similar distributions for $P$ and $\tau_{m}$.
Figure 21: Analysis of ring puckering motions. (A) Puckering occurs when atoms comprising the ribose sugar ring move above (endo) or below (exo) the ideal flat plane (cyan box). (B) Rose chart of phase angle $P$ distribution (left) and violin plot of amplitude $\tau_{m}$ distribution (right). RNA-FrameFlow (green) generates RNA backbones that capture an appropriate degree of puckering seen in natural RNA (purple).